Ana can paint a room in 9 1/2 hours.Karen can paint the sameroom in 6 1/2 hours.How lonh will it take them to finish painting the same room if they will work together?

Answer by richwmiller(17219)

a=9.5
b=6.5
1/a+1/b=1/x
1/9.5+1/6.5=1/x
1/9.5+1/6.5=1/3.859375
1/a+1/b=0.25910931
1/x=1/3.859375
x=3.859375